OpenStudy (anonymous):

Exactly 313.5 J will raise the temperature of 10.0 g of a metal from 25.0 C to 60.0 C. What is the specific heat of the metal?

OpenStudy (rogue):

\[Q = mc \Delta T\]\[313.5 J = 10.0 g \times 35°C \times c\]\[c = \frac {313.5 J}{350. g°C} = 0.896 J/(g°C)\]